OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

now. do u see the term 4n(4n+2) on the left side of the equation??

OpenStudy (anonymous):

Yes. That's actually what it says in the problem.

OpenStudy (helder_edwin):

this small formula is supposed to prepoduce all the termns on the left.

OpenStudy (helder_edwin):

if n=1 u get \[ \large 4\cdot1(4\cdot1+2)=4\cdot(4+2)=4\cdot6 \] which is fine. it is the first term. OK?

OpenStudy (helder_edwin):

@aleisha0301 ??

OpenStudy (anonymous):

Yes I understand so far

OpenStudy (helder_edwin):

the problem is that if u put n=2 u should get the second term but \[ \large 4\cdot2(4\cdot2+2)=8\cdot(8+2)=8\cdot10 \] which is not the second term !!!!

OpenStudy (anonymous):

So therefore the statement is false?

OpenStudy (helder_edwin):

it is either that or the problem is ill-posed. but let's write down that it is false.

OpenStudy (anonymous):

Yes that's the easier answer haha. But I have one last problem. Would you mind helping me solve it?

OpenStudy (helder_edwin):

not at all. but perhaps u could check the problem with one of your classmates. just to be sure.

OpenStudy (anonymous):

Unfortunately this is an online class :/

OpenStudy (helder_edwin):

oh. then let's try the other problem u mentioned.

OpenStudy (anonymous):

Alrighty